OpenStudy (anonymous):

Suppose g varies inversely as w, and when w is 8, g is 37. What is g when w is 5.2? Round your answer to two decimal places. Answer

OpenStudy (anonymous):

g=x/w is relationship methinks where x is a constant 37=x/8 x=296 g=296/5.2 g=56.923

OpenStudy (anonymous):